### v1

# AnimateAnywhere: Rouse the Background in Human Image Animation https://animateanywhere.github.io/

Abstract.
Human image animation aims to generate human videos of given characters and backgrounds that adhere to the desired pose sequence. However, existing methods focus more on human actions while neglecting the generation of background, which typically leads to static results or inharmonious movements. The community has explored camera pose-guided animation tasks, yet preparing the camera trajectory is impractical for most entertainment applications and ordinary users. As a remedy, we present an AnimateAnywhere framework, rousing the background in human image animation without requirements on camera trajectories. In particular, based on our key insight that the movement of the human body often reflects the motion of the background, we introduce a background motion learner (BML) to learn background motions from human pose sequences. To encourage the model to learn more accurate cross-frame correspondences, we further deploy an epipolar constraint on the 3D attention map. Specifically, the mask used to suppress geometrically unreasonable attention is carefully constructed by combining an epipolar mask and the current 3D attention map. Extensive experiments demonstrate that our AnimateAnywhere effectively learns the background motion from human pose sequences, achieving state-of-the-art performance in generating human animation results with vivid and realistic backgrounds. The source code and model will be available at https://github.com/liuxiaoyu1104/AnimateAnywhere.
1. Introduction
Human image animation has become a popular task focused on transforming source character images into realistic videos based on human pose sequences. This technique has a wide range of potential applications, including in entertainment videos, artistic creation, and virtual characters.
Recent advancements in generative diffusion models have significantly advanced the field of image animation (Hu, 2024; Xu et al., 2024b; Zhu et al., 2024; Zhou et al., 2024; Wang et al., 2024e, d), enabling the creation of highly realistic and controllable human image animation, particularly in terms of identity preservation and motion consistency. Despite these advancements, most existing human image animation methods focus on improving the generation quality of foreground humans while leaving the background static or unconstrained. However, in practice, the photographers usually need to adjust the cameras to better capture human movements, postures, and etc., which results in background changes. Thus, the static or disharmonious backgrounds are insufficient to meet the requirements for realistic and diverse human video generation.
To address this issue, one possible solution is to employ control signals to guide background motion. For example, Humanvid (Wang et al., 2024a) controls the background motion via camera trajectories estimated from the reference videos, while Liu et al. (Liu et al., 2024) leverage sparse tracking points and human pose sequences to independently control background and foreground motions. However, these methods typically acquire background control signals from a reference video, which is unavailable in many scenarios. Although camera trajectories can be manually designed, this easily results in disharmony between human and background motion. It requires professional skills and is redundant for most applications and ordinary users. Additionally, another limitation of these methods is that a reference image with both the human and the background is required.
To overcome the above limitations of existing human image animation methods, this work proposes AnimateAnywhere, a novel framework that generates videos with harmonious human and background motion controlled solely by human pose sequences, without additional camera control conditions. As shown in fig. 1, the key idea of our AnimateAnywhere is to identify the background motion from human pose sequences. We argue that human pose sequences can imply background motion. For instance, when a person appears larger or smaller across frames while remaining stationary, the background typically exhibits a corresponding zoom-in or zoom-out motion. Moreover, if a person walks to the left while remaining centered in the frame, it means that the camera is moving left, and the background should shift accordingly. By observing the relative changes in a person’s size and position across frames, as well as the type of human actions that can also be inferred from the human pose sequence, we can estimate both the direction and extent of background motion.
Specifically, we build the AnimateAnywhere framework upon a pre-trained image-to-video model (i.e., CogVideoX (Yang et al., 2024b)) and independently inject a reference human image, a reference background image, and a human pose sequence into the generation process. Thus, AnimateAnywhere supports free combinations of reference humans, backgrounds, and pose sequences, enabling highly flexible and controllable video generation. To realize the key idea of AnimateAnywhere, we propose the Background Motion Learner (BML) module, which learns background motion from the latent features of human pose sequences by applying LoRA (Hu et al., 2022) on the 3D-attention-based block. Moreover, although 3D full attention allows each background pixel to attend to any other pixel from all frames, this highly free attention manner also brings difficulties to optimization, easily resulting in geometrically inconsistent background generation. To further improve the cross-frame geometric consistency of the background, we propose an adaptive epipolar constraint for the 3D attention map. Therein, a mask used for suppressing geometrically unreasonable attention is elaborately designed from an epipolar mask and the current 3D attention map.
As shown in fig. 1, our AnimateAnywhere can generate realistic videos with harmonious human and background motion, without additional background motion control conditions. Extensive experiments conducted on the constructed datasets show that our AnimateAnywhere outperforms state-of-the-art methods, demonstrating the effectiveness, flexibility, and controllability of generating dynamically harmonious human and background motions. Our contributions are listed as follows.
We propose AnimateAnywhere, a novel human image animation framework that generates dynamic backgrounds in harmony with human motions, requiring no camera pose information. It also supports flexible combinations of reference humans, backgrounds, and pose sequences.
We propose a Background Motion Learner (BML) module to predict background motion from human pose sequences and an adaptive epipolar constraint for the 3D attention map to enhance background geometric consistency.
Our method achieves state-of-the-art performance in the challenging test dataset, demonstrating the effectiveness, flexibility, and controllability for human animation.
2. Related Work
2.1. Human Image Animation
Human image Animation aims to generate coherent human videos from a reference image and a target pose sequence. By leveraging the generative visual capabilities of diffusion models, recent advancements (Hu, 2024; Xu et al., 2024b; Zhu et al., 2024; Zhou et al., 2024; Wang et al., 2024e, d, a; Yao et al., 2025) have enabled the generation of high-quality and consistent human animation videos. AnimateAnyone (Hu, 2024) and MagicAnimate (Xu et al., 2024b) introduce a video diffusion model and a reference network for temporal consistency and identity preservation, respectively. Champ (Zhu et al., 2024) incorporates four distinct control signals to provide comprehensive 3D shape and detailed pose guidance. Following these, some works focus on realistic hands (Zhou et al., 2024; Wang et al., 2024e), efficient temporal modeling (Wang et al., 2024d), or a pose alignment algorithm (Tong et al., 2024) to further improve the performance of human animation. Animate-X (Tan et al., 2024) extends animation to diverse character types with a unified framework.
Despite achieving reasonable results in maintaining temporal consistency in humans, these approaches face challenges in generating realistic and dynamic backgrounds. MIMO (Men et al., 2024) and AnimateAnyone2 (Hu et al., 2025) address this challenge by taking the background video directly from a reference video as input. Liu et al. (Liu et al., 2024) extract sparse point tracking from the reference video and use it to guide background animation. Humanvid (Wang et al., 2024a) utilizes a SLAM-based method (Wang et al., 2024b; Teed and Deng, 2021) to estimate an accurate camera trajectory from the original videos and leverages CameraCtrl (He et al., 2024) and Animate Anyone (Hu, 2024) to construct a human animation framework. DynaScene (Yao et al., 2025) further introduces a multi-task learning approach to enhance the realism of the generated background . In comparison to the existing methods, our AnimateAnywhere predicts the background motion from human pose sequences without the reference video and camera pose extraction during inference.
2.2. Camera Control Video Generation
With the impressive capabilities of video diffusion models (Ho et al., 2022; Guo et al., 2023; Blattmann et al., 2023; Bar-Tal et al., 2024; Gupta et al., 2025) in generating high-quality videos, several approaches have focused on achieving controllable video generation by incorporating conditional controls (Zhang et al., 2023b; Guo et al., 2025; Wei et al., 2024; Guo et al., 2023; He et al., 2024). In this section, we provide a brief overview of the progress in camera-controllable video generation. AnimateDiff (Guo et al., 2023) uses MotionLoRA (Hu et al., 2021) fine-tuning to adapt a pre-trained video model to specific camera movements. Direct-a-video (Yang et al., 2024a) introduces a camera embedder that incorporates three camera parameters, though the limited parameters restrict camera movement to translations and zooms. Following this, MotionCtrl (Wang et al., 2024c) incorporates rotation and translation matrices derived from extrinsic camera parameters, enabling the generation of videos with more complex camera motions. CameraCtrl (He et al., 2024) leverages Plücker coordinates to represent camera poses, allowing for geometric interpretations at the pixel level. Recent advances further enhance the geometric consistency of generated videos by epipolar constraints or 3D reconstruction. CamCo (Xu et al., 2024a) applies epipolar constraints to ensure consistency between generated frames and a single reference frame only. CAMI2V (Zheng et al., 2024) extends epipolar constraints onto all frames to account for scenarios where frames lack overlapping regions with the reference frame. CamTrol (Hou et al., 2024) models camera movements through explicit 3D point clouds. Building on the epipolar constraint used in CAMI2V (Zheng et al., 2024), we introduce an adaptive epipolar constraint on the 3D attention map of the background regions.
3. Method
Given a human image , a background image , and a human pose sequence , AnimateAnywhere generates a video with synchronized human and background motion, guided solely by the pose sequence without explicit background motion control. Firstly, the preliminaries of AnimateAnywhere, including our base model CogVideoX (Yang et al., 2024b) and 3D Full Attention, are introduced in section 3.1. Additionally, we introduce the Background Motion Learner (BML) module that infers background motion representations from the latent features of the human pose sequence (as detailed in section 3.2). To further enhance the accuracy of background motion prediction, we incorporate epipolar masks derived from labeled camera trajectories as supervision (see section 3.3). Finally, the training objectives for AnimateAnywhere are described in section 3.4.
3.1. Preliminaries
Image-to-video CogVideoX Model. CogVideoX (Yang et al., 2024b) is a video diffusion model with transformers (DiTs) that achieves high-quality video generation. Our AnimateAnyWhere is built upon the Image-to-Video framework of CogVideoX (Yang et al., 2024b). Given a condition image, CogVideoX (Yang et al., 2024b) first pads it with zeros to N frames. Then the padded input is encoded by VAE and concatenated with noise. Subsequently, the combined latent is iteratively denoised over several steps to generate the final video. CogVideoX (Yang et al., 2024b) employs unified 3D full attention modules within each DiT block. The 3D full attention module performs calculations simultaneously in spatial and temporal dimensions, capturing both intra-frame and cross-frame interactions. This enables CogVideoX (Yang et al., 2024b) to comprehensively aggregate temporal information for video generation.
3D Full Attention. Firstly, the video latent is reshaped as , where , , and denote the temporal, channel, and spatial dimensions, respectively. And then the reshaped latent is input into the 3D Full Attention module. This 3D Full Attention module achieves a better understanding of video content by unifying the modeling of both time and space. The attention map is computed as:
where and denote the query and key projections, respectively. The resulting attention map reflects the correspondences across video frames learned by the diffusion model.
3.2. AnimateAnywhere
As shown in fig. 2, AnimateAnywhere aims to simultaneously animate both the reference human and the reference background. The animation of the human is guided by the human pose sequence, while the background is guided by the implicit motion signals derived from the human pose sequence. Below, we describe the detailed process of human and background animation.
Human Animation. (1) Reference Human Input. To ensure that the human details are well preserved during animation, we introduce a ReferenceNet DiT, whose structure matches the first 18 blocks of the denoising DiT. Specifically, the reference human image is first encoded into latent space using a 3D VAE and subsequently passed through the ReferenceNet DiT to extract human features . The human features are concatenated with the features from the main denoising DiT. These combined features are then processed through 3D full attention layers while preserving the original dimensions of the denoising DiT. (2) Human Pose Sequence Guidance. We adopt an architecture similar to ControlNet (Zhang et al., 2023a; Mou et al., 2024) to guide human motion. Specifically, we replicate the first 18 blocks of the pre-trained denoising DiT to serve as the ControlNet DiT. We utilize the VAE encoder to transform the human pose sequence into latent features and then utilize the ControlNet DiT to extract multi-scale pose features . Subsequently, these multi-scale features are added to the corresponding features of the denoising DiT via zero-initialized linear layers.
Background Animation. (1) Reference Background Input. Given a condition image, the CogVideoX (Yang et al., 2024b) I2V model uses it as the first frame to generate a video. Following this setup, we treat the reference background as the condition image and the initial background of the generated video. The background image is initially padded to match the dimension of the video, then encoded into latent features via the VAE encoder, and finally concatenated with noise to serve as input for the DiT module. (2) Background Motion Learner Module. Unlike other methods that extract camera trajectories from reference videos as background motion signals, we propose to directly predict background motion from human pose sequences. An intuitive approach is to learn camera trajectories from human pose sequences and incorporate the learned camera trajectories during video generation. However, directly learning camera trajectories faces several challenges. First, a single human pose may correspond to multiple plausible camera trajectories due to the ambiguity of whether the human is moving or stationary. Moreover, the labeled camera trajectories often suffer from scale inconsistency, which also increases the optimization difficulty. In contrast, rather than explicitly decoupling camera trajectories, we directly utilize the background motion features embedded in human pose to drive background movement. In particular, we introduce a LoRA-based module named Background Motion Learner (BML) module into each DiT block, and this module follows the standard paradigm of LoRA (Hu et al., 2022), expressed as . We add the human pose features and the denoising features together to predict the background motion. The process is as follows:
where and denote DiT block and BML module. Specifically, has strong priors on the co-movement between humans and backgrounds. , which is extracted from 3D attention in ControlNet, contains temporal information of human motion across frames.
3.3. Epipolar Constraint on 3D Attention Map
Although 3D full attention allows each background pixel to attend to any other pixel from all frames, this highly free search manner also makes optimization difficult and often causes geometric inconsistency in the background. To address this, we introduce epipolar constraints into the 3D attention map. However, strictly restricting the correspondence search space to geometrically reasonable regions may undermine the cross-frame correlations learned by the original 3D full attention. So we propose an adaptive epipolar constraint that integrates an epipolar mask and the current 3D attention map. In the following, we first introduce the preliminaries of the epipolar constraint and the details of our proposed adaptive epipolar constraint on the 3D attention map.
Epipolar Line and Epipolar Mask. The camera poses consist of intrinsic parameters and extrinsic parameters , where represents the rotational matrix and ddenotes the translational matrix. The epipolar constraint establishes that a point in the -th frame must lie on its corresponding epipolar line in the -th frame, as illustrated in fig. 3. The epipolar line can be determined by camera poses as:
where is the fundamental matrix between two frames. This matrix is computed as . Here, and are the relative rotation matrix and translation vector from -th frame to -th frame. Using the epipolar line, we convert each pixel in the -th frame into an epipolar mask . The epipolar mask significantly reduces the number of matching points from (frame size) to (where ).
Adaptive Epipolar Constraint. We define the vanilla 3D attention map as the attention intensity, indicating how much the pixel in the -th frame attends to each position in the -th frame. As described above, the epipolar mask strictly defines the valid region in the -th frame that the pixel in the -th frame should correspond to under projective geometry. To ensure that background motion adheres to geometric constraints, for each background pixel in -th frame, we suppress the attention values of that fall outside the corresponding epipolar region . However, completely suppressing attention outside the epipolar region may disrupt the cross-frame correlations learned by 3D full attention. To mitigate this, we adaptively apply the epipolar constraint by incorporating a dynamic attention mask defined as (), where () represents the region of low-confidence attention scores, i.e., the regions the network itself tends to ignore. Thus, the suppressed regions are determined by the intersection: . The final epipolar constraint is then defined by,
where denotes a statistical threshold corresponding to the 30th percentile of attention values. Meanwhile, we emphasize that our loss function is equivalent to computing the difference between and the target map, as illustrated in fig. 2 (b)-(5). During inference, our AnimateAnywhere no longer requires epipolar information or camera trajectories, and it directly leverages the attention learned through the epipolar loss.
3.4. Learning Objective
We employ the latent loss in CogVideoX (Yang et al., 2024b) to train our pipeline,
where denotes our diffusion model, represents the latent embeddings of an -frame image sequence, and corresponds to the noisy latent at diffusion timestep . The variable denotes the conditioning inputs provided to our model, specifically including , , and , while denotes the unscaled noise.
Meanwhile, we employ a one-step sampling strategy to transform the predicted latent at timestep to and we decode into the RGB video through the VAE decoder. We employ the VGG perceptual loss between the generated frames and the target frames during the last 30% timesteps as follow:
where is the mask sequence of the human hand region.
The overall learning objective for training is defined by,
where and are the weights to balance the loss terms.
4. Experiments
4.1. Experimental Details
Dataset. We collect 5K videos from the Humanvid (Wang et al., 2024a) dataset and 3K videos from the Internet for training, where the Humanvid (Wang et al., 2024a) dataset has the annotation of camera pose sequence. Our training dataset consists of videos captured under diverse camera conditions, including both static and moving cameras.
We evaluate our model on the Humanvid (Wang et al., 2024a) test dataset and a wild dataset, Bilibili200 (BL200), collected from the internet. The Humanvid test dataset includes 100 videos with noticeable background motion but limited human motion. The BL200 dataset contains 200 videos with significant motion in both the human and background.
For fair comparison, we evaluate all methods by generating 48-frame sequences with a stride of 4. Notably, while other methods utilize a single reference image that contains both the human and background during testing, our method inputs the reference human and reference background separately.
Evaluation Metrics. Following the metrics utilized in (Hu, 2024; Xu et al., 2024b; Zhu et al., 2024; Zhou et al., 2024; Wang et al., 2024e, d), we use PSNR (Huynh-Thu and Ghanbari, 2008), SSIM (Wang et al., 2004), and LIPIS (Zhang et al., 2018) metrics to evaluate the quality of single-frame images. Additionally, we introduce FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019) metrics to evaluate the difference between the distributions of generated videos and real-world videos.
Implementation Details. We adopt the pretrained CogVideoX1.5-5B-I2V (Yang et al., 2024b) as our base model and keep it frozen throughout the training process. Both ControlNet and ReferenceNet are initialized with the parameters of CogVideoX (Yang et al., 2024b). Our training consists of three stages. In the first stage, we incorporate ControlNet into the base model and train only ControlNet while keeping all other parameters fixed. In the second stage, we introduce ReferenceNet and train it exclusively. In the final stage, we jointly optimize ControlNet, ReferenceNet, and the BML module.
In terms of data preparation, we first employ Grounded-SAM (Ren et al., 2024) to extract mask sequences in the human body and hand regions for each video. Subsequently, we obtain the human sequence and the background sequence using the extracted mask sequence. We randomly select a frame in the human sequence as the reference human and take the first frame from the background sequence as the reference background. Specifically, we complete the reference background using the inpainting method (Zhou et al., 2023). All experiments are trained on a mixture of horizontal (1360768) and vertical (7681360) videos and conducted on two A100 GPUs. Training is performed over 20,000 steps with a batch size of 8. We employ the AdamW optimizer (Loshchilov and Hutter, 2017) with a constant learning rate of . We set the trade-off weights and to 0.2 and 0.005, respectively. During inference, classifier-free guidance (CFG) is applied with a scale of 2.5 for both reference images and pose sequences.
Methods | HumanVid | BL200 | ||||||||
PSNR | SSIM | LIPIS | FID | FVD | PSNR | SSIM | LIPIS | FID | FVD | |
MagicAnimate (Xu et al., 2024b) | 16.01 | 0.5463 | 0.4455 | 78.01 | 1943.07 | 16.82 | 0.5914 | 0.4133 | 40.02 | 1884.48 |
AnimateAnyone (Hu, 2024) | 16.47 | 0.5311 | 0.3913 | 50.18 | 1431.43 | 17.58 | 0.5756 | 0.3635 | 24.46 | 1507.05 |
Champ (Zhu et al., 2024) | 15.08 | 0.5075 | 0.4481 | 63.92 | 1837.07 | 16.88 | 0.5649 | 0.3919 | 30.69 | 1701.45 |
MusePose (Tong et al., 2024) | 16.93 | 0.5578 | 0.3676 | 52.00 | 1784.85 | 18.67 | 0.6126 | 0.3220 | 22.86 | 1580.47 |
AnimateAnyone† (Hu, 2024) | 16.81 | 0.5486 | 0.3672 | 47.69 | 1235.65 | 18.23 | 0.5930 | 0.3379 | 23.27 | 1371.22 |
MusePose† (Tong et al., 2024) | 17.33 | 0.5629 | 0.3563 | 44.73 | 1362.14 | 18.60 | 0.6116 | 0.3245 | 22.09 | 1444.79 |
Humanvid‡ (Wang et al., 2024a) | 16.42 | 0.5242 | 0.3676 | 47.41 | 1114.55 | - | - | - | - | - |
Ours | 18.07 | 0.5831 | 0.3119 | 44.51 | 978.09 | 18.99 | 0.6182 | 0.3206 | 21.31 | 1154.27 |
4.2. Comparison with Existing Methods
We compare our method with existing state-of-the-art human animation methods in both quantitative and qualitative aspects, including MagicAnimate (Xu et al., 2024b), AnimateAnyone (Hu, 2024), Champ (Zhu et al., 2024), MusePose (Tong et al., 2024), and HumanVid (Wang et al., 2024a). Since AnimateAnyone (Hu, 2024) is not officially open-sourced, we adopt its publicly available reimplementation from Moore-AnimateAnyone111https://github.com/MooreThreads/Moore-AnimateAnyone for evaluation. Note that these methods, except HumanVid (Wang et al., 2024a), are originally trained on datasets with static backgrounds. For fair comparison, we retrain AnimateAnyone (Hu, 2024) and MusePose (Tong et al., 2024) on our training dataset. In addition, Humanvid (Wang et al., 2024a) is capable of generating dynamic backgrounds, similar to ours. As its official implementation is not released, we reimplement Humanvid (Wang et al., 2024a) according to the details mentioned in its paper.
Quantitative Comparison. We conduct comprehensive experiments to assess the effectiveness of the proposed method, and the quantitative results are shown in table 1. From all metrics on the Humanvid (Wang et al., 2024a) test dataset and the BL200 test dataset, we can see that previous methods struggle to generate videos with dynamic backgrounds, resulting in the difference between the dynamic ground-truth videos. Even the retrained versions of these baselines occasionally produce static backgrounds, as they are not properly designed for background motion synchronized with human poses. In contrast, our method leverages CogVideoX (Yang et al., 2024b), the BML, and the epipolar loss to achieve substantial improvements, outperforming all baselines in reconstruction metrics. Furthermore, it also achieves state-of-the-art video fidelity, improving FID by 6% and FVD by 12% over the strongest baseline.
Qualitative Comparison. The qualitative comparison with competing methods is shown in fig. 4. MagicAnimate (Xu et al., 2024b), AnimateAnyone (Hu, 2024), Champ (Zhu et al., 2024), and MusePose (Tong et al., 2024) generate videos with static backgrounds that are usually unrealistic and exhibit various artifacts. Even after retraining, these methods show limited improvements and still fail in some circumstances. For instance, the left example exhibits the static background, while the right example produces an inharmonious scene in which only the wall moves and the table remains motionless. Humanvid is more likely to generate dynamic backgrounds, yet the background motion is frequently misaligned with the human. In the left example, the background moves noticeably slower than the person. In contrast, our method synthesizes photorealistic videos while maintaining coherent motion between the human and the background. fig. 5 illustrates more generated videos on BL200 dataset. These results also highlight our method’s capability to generate realistic background motion while maintaining high-quality, temporally consistent animation of both the human and the background. Moreover, our method achieves better identity-preserving than other methods.
PSNR | SSIM | LIPIS | FID | FVD | |
Ours w/o | 18.06 | 0.5807 | 0.3246 | 45.22 | 981.37 |
Ours with () | 17.32 | 0.5697 | 0.3280 | 50.81 | 1070.79 |
Ours with ()) | 18.07 | 0.5831 | 0.3119 | 44.51 | 978.09 |
Animation Cross Human and Background. Existing methods typically rely on a single reference image that includes both the human and the background, which limits the ability to customize videos for individual characters and backgrounds. In contrast, our method allows for separate input of the reference human and the reference background, enhancing the flexibility of the generation process. To validate the effectiveness of our approach, we present additional results of different characters and backgrounds in LABEL:{fig:Animate_cross}, further confirming the unique advantages of our method.
5. Ablation Study
Effects of Epipolar Loss. As shown in fig. 7, the generated video exhibits noticeable artifacts in the background without incorporating epipolar loss, because the learned background motion does not adhere to geometric constraints. Additionally, suppressing all attention values outside the epipolar mask does not alleviate this issue but performs even worse. As evidenced in the second row of fig. 7, noticeable artifacts and unexpected objects appear in the background. In contrast, our approach achieves desired results by applying to the region of . As shown in table 2, our method with outperforms both the one without epipolar loss and the one using the mask , demonstrating its effectiveness.
Effects of VGG Loss. We investigate the impact of the VGG perceptual loss on video generation. While the 3D VAE used in CogVideoX (Yang et al., 2024b) achieves efficient spatiotemporal compression (8 spatially, 4 temporally), we observe that optimizing solely with an L2 loss in latent space leads to degraded high-frequency details. As shown in fig. 8, the hand regions exhibit blurred contours without . After incorporating in pixel space, the details of the hand regions are well-preserved.
Effects of Foundational Models. To validate the superiority of our base model CogVideoX (Yang et al., 2024b), we select several state-of-the-art video generation base models (AnimateDiff (Guo et al., 2023), SVD (Blattmann et al., 2023)) for comparison. Specifically, we build two new pipelines based on AnimateDiff (Wang et al., 2024a) and SVD (Blattmann et al., 2023), and train them using our training dataset. Although these pipelines can generate videos with dynamic backgrounds, they struggle with the generation of novel regions. For instance, as shown in fig. 9, the pipeline based on AnimateDiff (Guo et al., 2023) generates blurred novel regions, while the pipeline based on SVD (Blattmann et al., 2023) yields unnatural human parts. In contrast, our method, based on CogvideoX (Yang et al., 2024b), successfully addresses these challenges.
6. Conclusion
In this paper, we introduce AnimateAnywhere, a human animation framework that generates videos with synchronized human and background motion without requiring specific background motion controls. The core of our framework is learning background motion representations directly from human pose sequences. Our method also provides additional geometric constraints on background motion learning by leveraging the labeled camera poses during training. Extensive experiments across multiple benchmarks validate the effectiveness of our AnimateAnywhere, demonstrating state-of-the-art performance in creating dynamic, realistic backgrounds that harmonize with human motion. This innovation not only improves the realism of the generated videos but also opens new possibilities in applications such as entertainment, virtual reality, and artistic content creation, where dynamic and immersive human-centered animations are increasingly valuable.